If a gas at a constant one atmosphere of pressure is heated from 27 C to 127 C what will be it's percentage increase in volume?
Answer #1

As PV = nRT

So

V = (nR/P)*T

And hence

V1 = (nR/P)*T1

V2 = (nR/P)*T2

dV = V2 - V1 = (nR/P)*(T2 - T1)

Hence fractional change = dV/V1 = (T2 - T1)/T1 = (400-300)/300 = 1/3 = 0.333 or 33.3%